5x-40=6y
2y=8-3x by the addition method

Answer by oberobic(2304)

The 'addition' method for solving systems of linear equations is also known as the 'elimination' method.
.
First, get the equations into standard form: Ax + By = C
.
5x -40 = 6y
5x - 6y = 40 :: (EQ 1)
.
2y = 8 - 3x
3x + 2y = 8 :: (EQ 2)
.
If we multiply EQ 2 by 3, the two equations will be in shape to eliminate the 'y' term.
.
5x -6y = 40 :: (EQ 1)
9x + 6y = 24 :: (EQ 2 * 3)
.
adding...
.
14x = 64
.
x = 64/14 = 32/7
.
substitute x = 32/7 in EQ 1 to find y
.
5(32/7) -6y = 40
160/7 -6y = 40
.
multiply both sides by 7 to remove the denominator
.
160 -42y = 280
.
subtract 120 from both sides
.
-42y = 120
.
divide both sides by -42
.
y = -120/42 = -60/21
.
So, we think the intersection of the two lines will be: (32/7, -60/21) = (4 4/7, -2 6/7)
.
Graphing is good way to check.
The slope-intercept form is a good way to compute the lines: y = mx+b.
.
5x-40=6y
.
6y = 5x -40
.
divide by 6
.
y = 5/6x - 40/6
y = 5/6x - 20/3
.
2y=8-3x
2y = -3x + 8
y = -3/2x + 4
.